OpenStudy (snitchseeker1496):

For the next two numbers, consider a uniform rod of length 5.41 m that is inclined at 33.70 with respect to the horizontal. The rod is moving at constant speed 0.458 c along the horizontal relative to a stationary observer. Here, c is the speed of light in vacuum. What is the value of the relativistic correction factor, gamma ?

OpenStudy (snitchseeker1496):

What will be length of the rod (in m) as measured by the stationary observer?

OpenStudy (ljetibo):

Did you try applying the formula for length contraction? $$L = \frac{L_0}{\sqrt{1-\frac{v^2}{c^2} }}$$